### v1

# Making the complete OpenAIRE citation graph easily accessible through compact data representation

Abstract
The OpenAIRE graph contains a large citation graph dataset, with over 200 million publications and over 2 billion citations. The current graph is available as a dump with metadata which uncompressed totals $\sim 2.5$ TB. This makes it hard to process on conventional computers. To make this network more available for the community we provide a processed OpenAIRE graph which is downscaled to 32 GB, while preserving the full graph structure. Apart from this we offer the processed data in very simple format, which allows further straightforward manipulation. We also provide a python pipeline, which can be used to process the next releases of the OpenAIRE graph.
Data paper
1 Overview
Repository location
We use two repositories, Zenodo to share our data and fixed version of the processing pipeline, and git repository on Codeberg to facilitate sharing and further development of the processing pipeline.
Data repository
Code repository
Context
The OpenAIRE graph (Rettberg and Schmidt, 2012; Manghi et al., 2025) is a large knowledge graph storing several kinds of research data. In this work we focus on extracting the citations and publications from that dataset, distilling it, and making it more accessible.
There are several other citations graphs; OpenAlex (Priem et al., 2022), Open Research Knowledge Graph (Jaradeh et al., 2019), Crossref (Hendricks et al., 2020) and OpenCitations (Peroni and Shotton, 2020) among others. These datasets differ in many ways, some datasets are only citation graphs, while others like OpenAlex and OpenAIRE are research knowledge graphs that record much more than just citations and publications. The datasets also differ in their publication coverage (Martín-Martín et al., 2021; Culbert et al., 2025) and how the works are indexed, e.g. how they determine taxonomy (Ciuciu-Kiss and Garijo, 2024). Details of publication coverage between OpenAIRE and other open datasets are not well explored and it is not sufficient to compare publication numbers since these datasets aggregate publications from multiple sources, thus a large node number might simply be due to subpar de-duplication.
The OpenAIRE initiative offer several APIs, a search engine and a complete dataset dump (Manghi et al., 2025). The API and search engine serve as accessible avenues for any researcher, however they are limited in the scope of data that can be efficiently accessed. If a researcher is interested in large portions of the graph, or the graph as a whole, then the dump is more suitable.
Processing the whole graph dump, require software coding skills, a large amount of memory, storage and computational power, something not readily accessible to many scholars in the humanities (Dederke et al., 2024). To fill this gap, we pre-process the dump, distilling it into a more managable size, and distributing it as CSVs, making the whole graph, more accessible. Similar work has been done for OpenAlex, to make that citation network more accessible (Caetano Machado Lopes and Chacko, 2024).
2 Method
The OpenAIRE graph consist of many entities and many relations. These are stored in many files in the dump, we only need the publication files (nodes) and the relation files (citations/edges) to extract the citation network. The publication files themselves contain a lot of information about publications, only some of which is relevant for most researchers. The relation files include many types of relations, for this dataset we are only interested in citations (the type ”Cites”). The following shows the generic steps we followed to extract the citation network and minimize memory footprint (the full code is available at [Github link]).
Steps
  1. 1.
    Download the complete OpenAIRE graph dump (Manghi et al., 2025), including all publication and relation files.
  2. 2.
    Extract relevant publication fields and flatten nested JSON structures to obtain a tabular representation.
  3. 3.
    Generate new more memory efficient(int32) nodeIDs, and build a hashtable for translating between OpenAIRE IDs and our nodeIDs.
  4. 4.
    Process relation files using PySpark (Zaharia et al., 2016), retaining only relations of type Cites by extracting source and target identifiers. Use hashtable to replace OpenAIRE IDs
  5. 5.
    Update all citation relations to reference the new identifiers.
This achieves a distilled dataset where the relations are efficiently stored using pairs of short integers.
Quality control
A set of validation scripts are run after processing to check for loss of data. The scripts check whether any entries from the original dump was lost and check for completeness of feature columns.
Output
The processing pipeline takes input of OpenAire full dump and transforms it into the files described below.
  • citations.csv – a simple edge list of the graph (all the citations). Each row has a simple form of two connected nodeIDs of e.g.:
    159486578,118392581
  • publications.csv – all graph nodes (publications), it contains only the nodeID and, when available, the DOI. Each row has a simple form of e.g.:
    14209,10.3931/e-rara-45685
  • publications_large.csv – includes the same number of nodes as publications.csv, but includes additional fields, e.g. title, authors, description, etc. For a full overview of fields, see Table 3.
  • pipeline.zip – processing pipeline which transform full OpenAIRE dump into the dataset files above. Full details is documented in included README.md. The source code is also present online at
    https://codeberg.org/Zmeos/OpenAIRE-citation-extraction.
Memory requirements
Below we summarize the hardware requirements for the transformed dataset and the original OpenAire full dataset dump:
Dataset Size on disk (GB) Memory usage (GB)
citations.csv $39$ $16$
publications.csv 66 $16$
publications_large.csv $179$ $309$
Full OA – edges $1820$ NA
Full OA – nodes $700$ NA
Table 1: Comparison of storage size and memory requirements of the full OpenAIRE (OA) dataset (release 2025-12-01) and its transformed versions. Memory usage corresponds to the amount of GB each dataset occupied when loaded into a pandas dataframe (pandas development team, 2020). Citations are loaded as int32.
3 Dataset Description
Repository name
Zenodo
Object name
An distilled OpenAIRE citation graph
Format names and versions
CSV
Creation dates
Based on the 2025-12-01 OpenAIRE dump.
Dataset creators
Joakim Skarding wrote the pipeline that processed the dump, Pavel Sanda supervised the work. The OpenAIRE initiative, of which this project is unaffiliated, created the OpenAIRE dump.
Language
English
License
CC BY 4.0
Publication date
2026-02-12
4 Reuse Potential
Citation networks are routinely used in a wide range of scientific fields. Among others this includes; research in historical trends of science (Frank et al., 2019; Drivas, 2024; Kitajima and Okamura, 2025; González-Márquez et al., 2024), sociology of scientific knowledge (Crothers et al., 2020; Carradore, 2022), network science (Costa and Frigori, 2024; Xiao et al., 2025) and as training sets for graph neural networks (Kipf, 2016; Leskovec and Sosič, 2016). Since the dataset is an evolving network, it can also be used to train temporal models, such as dynamic graph neural networks (Skarding et al., 2021).
Funding Statement
This work has been funded by a grant from the Programme Johannes Amos Comenius under the Ministry of Education, Youth and Sports of the Czech Republic, CZ.02.01.01/00/23_025/0008711.
Competing interests
The authors have no competing interests to declare.
Supplementary information
Example of reduction of a relation (citation/edge)
Our dump has the benefit of only focusing on citations, thus we do not need to indicate the type of citation. We only retain the information of the source ID and target ID. These IDs are also optimized to take up as little space as possible. The following is an example of how an edge is represented in the OpenAIRE JSON dump:
  "provenance": {
    "provenance": "Inferred by OpenAIRE",
    "trust": "0.9"
  "relType": {
    "name": "Cites",
    "type": "citation"
  "source": "doi_________::7e8d84fc096936557defb78d22cca97c",
  "sourceType": "product",
  "target": "dedup_wf_002::27d83ddfd6e54378d88445aa793d5cb8",
  "targetType": "product",
  "validated": false
An edge in our dump is simply:
159486578,118392581
Estimates of OpenAIRE JSON files uncompressed
Table LABEL:tab:estimate shows a rough estimate of the full uncompressed size of the OpenAIRE dump. The dump would require several orders of magnitude more RAM than most consumer computers.
Table 2: Estimate of OpenAIRE JSON files when uncompressed
Relation (edges) Size Estimate
One relation JSON (uncompressed) 6.5 GB
One relation JSON (compressed) 0.5 GB
Compression ratio $6.5/0.5=13$
Compressed relation folder size 10 GB
Estimated uncompressed per folder $13\times 10=130\,\text{GB}$
Number of relation folders 14
Total uncompressed (relations) $14\times 130=\textbf{1.82\,TB}$
Publication (nodes) Size Estimate
One publication JSON (uncompressed) 0.5 GB
One publication JSON (compressed) 0.1 GB
Compression ratio $0.5/0.1=5$
Compressed publication folder size 10 GB
Estimated uncompressed per folder $5\times 10=50\,\text{GB}$
Number of publication folders 14
Total uncompressed (publications) $14\times 50=\textbf{0.70\,TB}$
Combined Uncompressed Estimate 2.5 TB
Size of datafields in publications_large.csv
Field Explanation Memory usage (GB)
nodeId Unique internal identifier for the node (publication) 22
openaireId Identifier assigned by the OpenAIRE platform $18$
doi Digital Object Identifier of the publication $13$
title Title of the publication $28$
authors List of authors associated with the publication 2020
description Abstract or short description of the publication $192$
date Date when the publication was published $11$
container Journal, conference, or repository where it was published $13$
citations Number of times the publication has been cited 22
language Language in which the publication is written $10$
Table 3: Description of data fields of the large version of the dataset. The publications.csv only contains the nodeId, doi, and citations. The memory usage is the amount of GB that column took when loaded into memory as a pandas dataframe.

### v3

@@
 The OpenAIRE graph contains a large citation graph dataset, with over 200 million publications and over 2 billion citations.
-The current graph is available as a dump with metadata which uncompressed totals $\sim 2.5$ TB.
+The current graph is available as a dump with metadata which, when uncompressed, totals $\sim 2.5$ TB.
 This makes it hard to process on conventional computers.
-To make this network more available for the community we provide a processed OpenAIRE graph which is downscaled to 32 GB, while preserving the full graph structure.
-Apart from this we offer the processed data in very simple format, which allows further straightforward manipulation.
-We also provide a python pipeline, which can be used to process the next releases of the OpenAIRE graph.
+To make this network more accessible for the community, we provide a processed OpenAIRE graph which is downscaled to 16 GB RAM, while preserving the full graph structure.
+Apart from this we offer the processed data in a very simple format, which allows for further straightforward manipulation.
+We also provide (1) a Python pipeline, which can be used to process the next releases of the OpenAIRE graph, and (2) a larger version of the dataset including more publication fields such as, the title, list of authors.
 Data paper
+Keywords: citation network; dynamic network; OpenAIRE; large scale network
 1 Overview
@@
 Context
-The OpenAIRE graph (Rettberg and Schmidt, 2012; Manghi et al., 2025) is a large knowledge graph storing several kinds of research data.
-In this work we focus on extracting the citations and publications from that dataset, distilling it, and making it more accessible.
-There are several other citations graphs; OpenAlex (Priem et al., 2022), Open Research Knowledge Graph (Jaradeh et al., 2019), Crossref (Hendricks et al., 2020) and OpenCitations (Peroni and Shotton, 2020) among others.
-These datasets differ in many ways, some datasets are only citation graphs, while others like OpenAlex and OpenAIRE are research knowledge graphs that record much more than just citations and publications.
+The OpenAIRE graph (Rettberg & Schmidt, 2012; Manghi et al., 2025) is a large knowledge graph storing several kinds of research data.
+In this work we focus on extracting the citations and publications from that dataset, distilling the information to make it more accessible.
+There are several other citations graphs available – OpenAlex (Priem et al., 2022), Open Research Knowledge Graph (Jaradeh et al., 2019), Crossref (Hendricks et al., 2020) and OpenCitations (Peroni & Shotton, 2020).
+These datasets differ in many ways, some are only citation graphs, while others like OpenAlex and OpenAIRE are research knowledge graphs that record much more than just citations and publications.
 The datasets also differ in their publication coverage (Martín-Martín et al., 2021; Culbert et al., 2025) and how the works are indexed, e.g.
-how they determine taxonomy (Ciuciu-Kiss and Garijo, 2024).
-Details of publication coverage between OpenAIRE and other open datasets are not well explored and it is not sufficient to compare publication numbers since these datasets aggregate publications from multiple sources, thus a large node number might simply be due to subpar de-duplication.
-The OpenAIRE initiative offer several APIs, a search engine and a complete dataset dump (Manghi et al., 2025).
-The API and search engine serve as accessible avenues for any researcher, however they are limited in the scope of data that can be efficiently accessed.
+how they determine taxonomy (Ciuciu-Kiss and Garijo, 2024b).
+Details of publication coverage between OpenAIRE and other open datasets are not well explored – a subset of OpenAIRE has been compared to OpenAlex (Ciuciu-Kiss and Garijo, 2024a), but to the best of our knowledge, no coverage analysis to the extent of previous works like (Martín-Martín et al., 2021) or (Culbert et al., 2025) has taken place.
+It is not sufficient to compare publication numbers since these datasets aggregate publications from multiple sources and a large node number might simply be due to poor de-duplication.
+The OpenAIRE initiative offers several APIs, a search engine, cloud access through Google BigQuery, and a complete dataset dump (Manghi et al., 2025).
+The API and search engine serve as accessible avenues for any researcher, however, they are limited in the scope of data that can be efficiently accessed.
+A scan of the complete graph is outside the free tier of BigQuery, at the time of writing.
 If a researcher is interested in large portions of the graph, or the graph as a whole, then the dump is more suitable.
-Processing the whole graph dump, require software coding skills, a large amount of memory, storage and computational power, something not readily accessible to many scholars in the humanities (Dederke et al., 2024).
-To fill this gap, we pre-process the dump, distilling it into a more managable size, and distributing it as CSVs, making the whole graph, more accessible.
-Similar work has been done for OpenAlex, to make that citation network more accessible (Caetano Machado Lopes and Chacko, 2024).
+Processing the whole graph dump requires software coding skills, a large amount of memory, storage and computational power, resources not readily accessible to many scholars in the humanities (Dederke et al., 2024).
+Accessing the raw data also requires solid understanding of the OpenAIRE JSON schema.
+To fill this gap, we pre-process the dump, distilling it into a more manageable size, and distributing it as simply structured plain text and Parquet files, making the whole graph, more accessible.
+Similar work has also been done for OpenAlex (Caetano Machado Lopes and Chacko, 2024).
 2 Method
-The OpenAIRE graph consist of many entities and many relations.
-These are stored in many files in the dump, we only need the publication files (nodes) and the relation files (citations/edges) to extract the citation network.
-The publication files themselves contain a lot of information about publications, only some of which is relevant for most researchers.
-The relation files include many types of relations, for this dataset we are only interested in citations (the type ”Cites”).
-The following shows the generic steps we followed to extract the citation network and minimize memory footprint (the full code is available at [Github link]).
+The OpenAIRE graph consists of many entities and many relations.
+These are stored in many files in the dump.
+We only need the publication files (nodes) and the relation files (citations/edges) to extract the citation network.
+The publication files themselves contain a lot of information about publications, only some of which is relevant to most researchers.
+The relation files include many types of relations.
+For this dataset we are only interested in citations (the type ”Cites”).
+The following list shows the generic steps we followed to extract the citation network and minimize memory footprint.
+The list is organized by python files and shows the file responsible for each step.
+(the full code is available in the Codeberg repository).
 Steps
+  1. 0.
+    step0_download_data_and_extract.py
+    Download the complete OpenAIRE graph dump (Manghi et al., 2025), including all publication and relation files. When calling the pipeline, this is an optional step.
   2. 1.
-    Download the complete OpenAIRE graph dump (Manghi et al., 2025), including all publication and relation files.
+    step1_extract_raw.py
+    Partially extract the compressed dump files.
   3. 2.
-    Extract relevant publication fields and flatten nested JSON structures to obtain a tabular representation.
+    step2_publications.py
+    1. (a)
+      Filters records, obtaining publication-type only.
+    2. (b)
+      Flatten nested JSON structures to obtain a tabular representation.
+    3. (c)
+      Generate new, more memory-efficient(int32) nodeIDs, and build a hashtable for translating between OpenAIRE IDs and our nodeIDs.
+    4. (d)
+      For the TSV files; replace any tabs or newlines with whitespace
+    5. (e)
+      Export TSV and Parquet publication files
   4. 3.
-    Generate new more memory efficient(int32) nodeIDs, and build a hashtable for translating between OpenAIRE IDs and our nodeIDs.
+    step3_citations.py
+    1. (a)
+      Process relation files using PySpark (Zaharia et al., 2016), retaining only relations of type Cites and extracting source and target identifiers.
+    2. (b)
+      Use hashtable produced by step2_publications.py to replace OpenAIRE IDs. Update all citation relations to reference the new identifiers.
+    3. (c)
+      Export TSV and Parquet citation files
   5. 4.
-    Process relation files using PySpark (Zaharia et al., 2016), retaining only relations of type Cites by extracting source and target identifiers. Use hashtable to replace OpenAIRE IDs
-  5. 5.
-    Update all citation relations to reference the new identifiers.
+    step4_distribution.py
+    Compresses the TSV files to xz.
 This achieves a distilled dataset where the relations are efficiently stored using pairs of short integers.
 Quality control
-A set of validation scripts are run after processing to check for loss of data. The scripts check whether any entries from the original dump was lost and check for completeness of feature columns.
+Several validation scripts are run after processing, to control for mistakes in the processing. The scripts check whether any entries from the original dump were lost and report any missing values.
+Script Metric Value
+full_id_coverage (pass)
+raw_count 205841448
+parquet_count 205841448
+missing_count 0
+extra_count 0
+distinct_constraints (pass)
+publications total 205841448
+publications distinct 205841448
+publications duplicates 0
+citations total 2184347684
+citations distinct 2184347684
+citations duplicates 0
+citations null_rows 0
+format_checks (pass)
+nodeid gaps 0
+nodeid contiguous true
+Table 1: Combined validation script output. This is the automated output log of the validation scripts.
+  • full_id_coverage.py Verifies that every publication in the raw source data is present in the final output.
+  • distinct_constraints.py Checks that there are no duplicate publications or duplicate citation edges in the output.
+  • format_checks.py Verifies that publication node IDs form a complete, gap-free sequence starting from zero.
+Additionally the file run_all_validations.py runs all validation scripts, collects their results, and writes a consolidated summary report.
+The entire validation pipeline is automatically run after the processing pipeline completes.
 Output
-The processing pipeline takes input of OpenAire full dump and transforms it into the files described below.
-  • citations.csv – a simple edge list of the graph (all the citations). Each row has a simple form of two connected nodeIDs of e.g.:
-    159486578,118392581
-  • publications.csv – all graph nodes (publications), it contains only the nodeID and, when available, the DOI. Each row has a simple form of e.g.:
-    14209,10.3931/e-rara-45685
-  • publications_large.csv – includes the same number of nodes as publications.csv, but includes additional fields, e.g. title, authors, description, etc. For a full overview of fields, see Table 3.
-  • pipeline.zip – processing pipeline which transform full OpenAIRE dump into the dataset files above. Full details is documented in included README.md. The source code is also present online at
+The processing pipeline takes the full OpenAIRE dump as input and transforms it into the TSV (tab-separated values) and equivalent Parquet files described below.
+TSV is a variant of a more common CSV.
+We chose TSV because several of the text fields in the publications_large.tsv routinely contained commas, and our attempts at quoting the commas turned out to be error prone.
+Parquet is a binary format for high-performance data processing.
+  • citations.tsv.xz and citations.parquet – a simple edge list of the graph (all the citations). Each row has a simple form of two connected nodeIDs, e.g.: 159486578 118392581
+  • publications.tsv.xz and publications.parquet – all graph nodes (publications), it contains only the nodeID and, when available, the DOI.
+Each row has a simple form, e.g.: 14209 10.3931/e-rara-45685
+  • publications_large.tsv.xz and publications_large.parquet– includes the same number of nodes as the publication files, but includes additional fields, e.g.
+title, authors, description, etc.
+For a full overview of fields, see Table 3.
+  • pipeline.tar.xz – processing pipeline which transform full OpenAIRE dump into the dataset files above. Full details are documented in included README.md. The source code is also present online at
     https://codeberg.org/Zmeos/OpenAIRE-citation-extraction.
+The pid_* columns (included in publications_large, see Table 3) are only extracted schemes from the pids field in the dump which OpenAIRE populates with identifiers collected from authoritative source.
+The instances field (a nested field of additional metadata, Table 5) also includes identifiers, those were not collected, with the exception of the primary_doi field in the publications file, there the doi is drawn from the identifiers field if it is not found in the pids field.
 Memory requirements
-Below we summarize the hardware requirements for the transformed dataset and the original OpenAire full dataset dump:
+Below, we summarize the hardware requirements for the transformed dataset and the original OpenAire full dataset dump.
+For loading the citations.tsv into memory with int32 using the Pandas library in Python, the following can be used:
+df_refs = pd.read_csv(
+[…]
+dtype={"source_nodeId": "int32", "target_nodeId": "int32"}
+The data formats (TSV and Parquet) are of course generic and can be loaded using any tool that supports these common formats. When loading the Parquet, int32 will be used automatically.
+df_cites = pd.read_parquet(
+"citations.parquet",
+engine="pyarrow",
+dtype_backend="pyarrow",
 Dataset Size on disk (GB) Memory usage (GB)
-citations.csv $39$ $16$
-publications.csv 66 $16$
-publications_large.csv $179$ $309$
+citations.tsv $39$ $17$
+publications.tsv 66 55
+publications_large.tsv $187$ $185$
+citations.parquet 88 $17$
+publications.parquet 22 55
+publications_large.parquet $68$ $185$
 Full OA – edges $1820$ NA
 Full OA – nodes $700$ NA
-Table 1: Comparison of storage size and memory requirements of the full OpenAIRE (OA) dataset (release 2025-12-01) and its transformed versions.
-Memory usage corresponds to the amount of GB each dataset occupied when loaded into a pandas dataframe (pandas development team, 2020).
-Citations are loaded as int32.
+Table 2: Comparison of storage size and memory requirements of the full OpenAIRE (OA) dataset (release 2025-12-01) and its compact versions.
+Memory usage corresponds to the amount of GB each dataset occupied when loaded into a Pandas dataframe (The pandas development team, 2026).
+Since the citations are loaded as int32, the memory size is much lower than disk size.
+The publications files benefit from being loaded using the PyArrow backend. This significantly reduces the memory usage of the doi field.
+df_pubs = pd.read_parquet(
+"publications.parquet",
+engine="pyarrow",
+dtype_backend="pyarrow",
+The publications_large files are the most important to load efficiently as $\sim 200$GB of RAM can be saved.
+Be aware that even efficient loading still requires $\sim 185$GB of RAM.
+For users interested in a subset of the provided columns, we recommend selecting the columns on load.
+For loading the full dataset the following code can be used, and selected columns can be removed.
+This efficient loading is the approach used for the ”Pandas Opt” column in Table 3.
+In the example below, the columns nodeId, title and pid_dois are selected.
+df_large = pd.read_parquet(
+"publications_large.parquet",
+columns=["nodeId", "title", "pid_dois"],
+engine="pyarrow",
+dtype_backend="pyarrow",
+df_large[["language", "container"]] = (
+df_large[["language", "container"]].astype("category")
+Column Python Type Description Arrow Pandas Pandas Opt
+nodeId int32 Node ID 0.767 GB 0.767 GB 0.767 GB
+openaireId str OpenAIRE unique ID 9.585 GB 19.746 GB 9.586 GB
+title str Paper title 16.486 GB 29.707 GB 16.486 GB
+authors list[str] List of authors 11.037 GB 23.005 GB 11.038 GB
+description str Main text/abstract 131.248 GB 193.583 GB 131.248 GB
+date datetime Publication date 0.767 GB 7.587 GB 0.768 GB
+container str Journal/conference name 5.471 GB 13.953 GB 2.181 GB
+citations int Citation count 1.558 GB 1.534 GB 1.558 GB
+language str Language 1.394 GB 11.555 GB 0.197 GB
+pid_dois list[str] DOI identifiers 5.639 GB 19.453 GB 5.639 GB
+pid_mag_ids list[str] MAG IDs 2.004 GB 12.748 GB 2.005 GB
+pid_pmids list[str] PubMed IDs 1.202 GB 7.948 GB 1.202 GB
+pid_handles list[str] Persistent handles 1.149 GB 6.134 GB 1.149 GB
+pid_pmcs list[str] PubMed Central IDs 0.921 GB 5.480 GB 0.921 GB
+pid_arxiv_ids list[str] ArXiv IDs 0.885 GB 4.856 GB 0.886 GB
+TOTAL 190.113 GB 358.055 GB 185.631 GB
+Table 3: Memory size of columns with Python types and short descriptions (Arrow, Pandas, and optimized Pandas formats).
+Arrow is the memory size loaded using PyArrow, Pandas is the size if the data is loaded straight into a default Pandas dataframe.
+The Pandas Optimized column uses PyArrow as a backend, and utilizes ”categorical” on the container and language fields to further lower their footprint.
+The MAG IDs are Microsoft Academic Graph IDs.
 3 Dataset Description
@@
 Object name
-An distilled OpenAIRE citation graph
+Compact representation of the OpenAIRE citation graph
 Format names and versions
-CSV
+TSV (Tab Separated Values) and Apache Parquet
 Creation dates
@@
 Dataset creators
-Joakim Skarding wrote the pipeline that processed the dump, Pavel Sanda supervised the work. The OpenAIRE initiative, of which this project is unaffiliated, created the OpenAIRE dump.
+Joakim Skarding wrote the pipeline that processed the dump, Pavel Sanda supervised the work. The OpenAIRE initiative, with which this project is unaffiliated, created the OpenAIRE dump.
 Language
@@
 Citation networks are routinely used in a wide range of scientific fields.
-Among others this includes; research in historical trends of science (Frank et al., 2019; Drivas, 2024; Kitajima and Okamura, 2025; González-Márquez et al., 2024), sociology of scientific knowledge (Crothers et al., 2020; Carradore, 2022), network science (Costa and Frigori, 2024; Xiao et al., 2025) and as training sets for graph neural networks (Kipf, 2016; Leskovec and Sosič, 2016).
+This includes among others, research in historical trends of science (Frank et al., 2019; Drivas, 2024; Kitajima & Okamura, 2025; González-Márquez et al., 2024), sociology of scientific knowledge (Crothers et al., 2020; Carradore, 2022), network science (Costa & Frigori, 2024; Xiao et al., 2025) and as training sets for graph neural networks (Kipf, 2016; Leskovec & Sosič, 2016).
 Since the dataset is an evolving network, it can also be used to train temporal models, such as dynamic graph neural networks (Skarding et al., 2021).
+Compared to working directly with the original OpenAIRE data dump, this dataset significantly reduces the time and effort required to begin analysis.
+The data is provided as flat TSV and Parquet files, removing the need to parse and process the raw JSON.
+The citation graph is already structured as an edge list with integer node IDs, the format expected by most graph libraries, meaning graph algorithms (e.g.
+community detection (Fortunato, 2010) and centrality measures (Bloch et al., 2023)) can be applied directly without further transformation.
+The reduced file size allows the full dataset to be downloaded and explored locally.
+Together, these properties make the dataset a convenient starting point for bibliometric studies, network analysis, and graph learning research.
+We provide the source code used to produce the data, allowing researchers to run the pipeline when new versions of the OpenAIRE graph is released, as well as make their own customized distilled dataset with the fields they desire.
 Funding Statement
@@
 Example of reduction of a relation (citation/edge)
-Our dump has the benefit of only focusing on citations, thus we do not need to indicate the type of citation.
-We only retain the information of the source ID and target ID.
+Our dump has the benefit of focusing solely on citations, and thus we do not need to indicate the citation type.
+We only retain the source and target IDs.
 These IDs are also optimized to take up as little space as possible.
@@
 An edge in our dump is simply:
-159486578,118392581
+159486578   118392581
 Estimates of OpenAIRE JSON files uncompressed
@@
 Combined Uncompressed Estimate 2.5 TB
-Size of datafields in publications_large.csv
-Field Explanation Memory usage (GB)
-nodeId Unique internal identifier for the node (publication) 22
-openaireId Identifier assigned by the OpenAIRE platform $18$
-doi Digital Object Identifier of the publication $13$
-title Title of the publication $28$
-authors List of authors associated with the publication 2020
-description Abstract or short description of the publication $192$
-date Date when the publication was published $11$
-container Journal, conference, or repository where it was published $13$
-citations Number of times the publication has been cited 22
-language Language in which the publication is written $10$
-Table 3: Description of data fields of the large version of the dataset.
-The publications.csv only contains the nodeId, doi, and citations.
-The memory usage is the amount of GB that column took when loaded into memory as a pandas dataframe.
+Filtered fields
+We remove a lot of fields from the original OpenAIRE graph dump.
+Table 5 is an overview of the fields we have filtered out of the publications (nodes).
+Table 6 is an overview of the fields filtered out of the relations (edges).
+Field Description
+Bibliographic
+subTitle Alternative or explanatory title
+publisher Publishing entity
+version Version of the result
+sources Dublin Core dc:source
+contributors Contributing persons or entities
+subjects Keywords with scheme and provenance
+formats File formats
+coverages Coverage information
+originalIds Identifiers at original sources
+dateOfCollection When OpenAIRE last collected the record
+lastUpdateTimeStamp Timestamp of last update in OpenAIRE
+embargoEndDate Date embargo ends
+Open Access
+bestAccessRight Openest access right (code, label, scheme)
+isGreen True if green Open Access
+isInDiamondJournal True if published in a Diamond Journal
+openAccessColor gold, hybrid, or bronze
+publiclyFunded True if outcome of a funded project
+Impact indicators
+indicators.citationImpact.citationClass Citation class label
+indicators.citationImpact.impulse Impulse score
+indicators.citationImpact.impulseClass Impulse class label
+indicators.citationImpact.influence Influence score
+indicators.citationImpact.influenceClass Influence class label
+indicators.citationImpact.popularity Popularity score
+indicators.citationImpact.popularityClass Popularity class label
+indicators.usageCounts.downloads Download count
+indicators.usageCounts.views View count
+Instances (per-version metadata)
+instances Per-version access rights, URLs, license, refereed status
+Geolocation
+geoLocations Box, place, and point geolocation data
+countries Associated countries with ISO codes
+Software/dataset specific
+codeRepositoryUrl URL to source code repository
+contactGroups Groups responsible for the software
+contactPeople Persons responsible for the software
+documentationUrls URLs to software documentation
+programmingLanguage Programming language of software
+size Declared size of dataset
+tools Tools for interpretation of result
+Sub-fields of included objects (partially used)
+authors.name Author first name
+authors.surname Author surname
+authors.rank Author position
+authors.pid Author ORCID identifier
+container.edition Journal or proceeding edition
+container.ep End page
+container.iss Journal issue number
+container.issnLinking Linking ISSN
+container.issnOnline Online ISSN
+container.issnPrinted Printed ISSN
+container.sp Start page
+container.vol Volume
+container.conferenceDate Conference date
+container.conferencePlace Conference location
+language.label Language label in English
+Table 5: Fields present in the OpenAIRE product schema that are not included in the output dataset.
+Field Description
+Core relation fields
+sourceType Type of the source entity (e.g. product, project, organization)
+targetType Type of the target entity
+relClass High-level category of the relation (e.g. citation, affiliation)
+relType Specific relation type describing the semantic meaning
+Provenance and validation
+provenance Information about how the relation was generated or provided
+validated Indicates whether the relation has been validated
+validationDate Date when the relation was validated
+Additional metadata
+confidence Confidence score for the relation
+inference Indicates whether the relation was inferred or explicitly provided
+Table 6: Fields present in the OpenAIRE relation schema that are not included in the output dataset.
+The core relation field ”relType” is used for filtering to only obtain ”Cites” relations.
+And removal of dangling edges ensures that the output dataset only includes edges with sourceType and targetType ”product”.